OpenStudy (anonymous):

Can someone walk me thru the steps to solve this? 0.22y-0.6=0.12y+3-0.8y

OpenStudy (anonymous):

First, group like terms on one side, 0.22y - 0.12y + 0.8y = 3 + 0.6 Perform addition and subtraction on the figures 0.9y = 3.6 Divide both sides by 0.9 to allow y to stand alone y = 3.6 /0.9 y = 4

OpenStudy (anonymous):

I have to tell u in person i will take me to long to write down

OpenStudy (anonymous):

thanks... I have a few questions.. U wanna help me with a few?

OpenStudy (anonymous):

1/4x - 1/8x = 3 - 1/16x